OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250? I know n=8, a1=10, an=781,250. I just don't understand how to find the common ratio if a list of the terms is not given and the problem doesn't say what the common ratio is?

OpenStudy (campbell_st):

a term in a geometric sequence is found using \[T _{n} = ar^{n -1}\] in your question you have the 8th term, the 1st term and n \[781250 = 10 \times r^{8-1}\] then \[r^7 = 78125\] \[r = \sqrt[7]{78125} = 5\] now you can find the sum

OpenStudy (anonymous):

So the ratio is 5?